The time between customers arriving at a book store is exponentially distributed with a mean of 5 minutes.

(a) What is the probability that no customers arrive during a 30-minute period? (5)

(b) If the store opens at 9:00, what is the probability that the first customer of the day arrives between 9:10 and 9:20? (6)

(c) If the store opens at 9:00, what is the probability that the first customer arrives at exactly 9:15?


1
Expert's answer
2021-07-25T09:55:21-0400

Suppose a random variable X denotes the time(in minutes) between customer's arrivals.

Therefore:

"X\approx Exp\left(\lambda =\frac{1}{5}=0.2\right)"

"f\left(x\right)=\lambda e^{-\lambda x}\:"

"f\left(x\right)\:=0.2e^{-0.2x}\:for\:x\ge 0"


(a). No customers arrive during a 30 - minute period if the the time between customer's arrivals is greater than 30 minutes.

Required probability is given by


"P\left(X>30\right)=\int _{30}^{\infty }" "0.2e^{-0.2x}dx\:"


"=\left[-e^{-0.2x}\:\right]^{^{\infty }}_{30}"


"=\left[e^{-0.2x}\:\right]^{^{30\:}}_{\infty }"


"=e^{-0.2\times 30}-e^{-\infty }"


"=e^{-6}-0\:"


"=0.0025"

Hence, "0.0025" is the probability that no customers arrive during a 30 - minute period.


(b). Considering 9:00 as the origin of time(i.e. time zero), we obtain 9:10 as 10 minutes and 9:20 as 20 minutes . Required probability is given by;


"P\left(10<X<20\right)=\int _{10}^{20}0.2e^{-0.2x}\:dx"


"=\left[-e^{-0.2x}\right]^{^{20}}_{10}"



"=\left[e^{-0.2x}\right]^{^{10}}_{20}"


"=e^{-0.2\times 10}-e^{-0.2\times 20}"


"=e^{-2}-e^{-4}\:"


"=0.1353353-0.0183156"


"=0.1170"

Hence, "0.1170" is the probability that the first customer arrives between 9:10 and 9:20.


(c). Considering 9:00 as the origin of time (i.e. time zero), we obtain 9:15 as 15 minutes.

Required probability is given by;

"P\left(X=15\right)=\int _{15}^{15}0.2e^{-0.2x}dx"


"=\left[e^{-0.2x}\right]^{^{15}}_{15}"


"=e^{-0.2\times 15}-e^{-0.2\times 25}"


"=e^{-3}-e^{-3}"


"=0.04978707-0.04978707"


"=0"

Hence, is the probability that the first customer arrives at exactly 9:15